OpenStudy (anonymous):

What is the slope of the line through the points (2, 6) and (3, 5)? m = _?_

OpenStudy (anonymous):

2 - 3 over 6 - 5 = -1/-1, or 1

OpenStudy (amistre64):

slope is defined as the change in y with respect to the change in x subtract the point to find the change (2, 6) -(3, 5) ------ -1,1 ; and slope = y/x = -1

OpenStudy (amistre64):

merllet typos almost as bad as me ;)

OpenStudy (anonymous):

right one, down one, slope is -1